What is the solution to this equation
9f+16=70

Answer by Maths68(1474)

9f+16=70
Subtract 16 from both sides of the equation
9f+16-16=70-16
9f=54
Divide by 9 both sides of the equation
9f/9=54/9
f=6
Solution for f is 6